OpenStudy (anonymous):

find the magnitude: || 6-2i ||

OpenStudy (anonymous):

\[\left| 6.325 \right| \] arg(-19.435 deg)

OpenStudy (maheshmeghwal9):

\[\huge{z= x+iy.}\]Then\[\huge{|z|= \sqrt{x^2+y^2}.}\]

OpenStudy (maheshmeghwal9):

So; ||6-2i||=\[|\sqrt{6^2+(-2^2)}|=|\sqrt{40}|\]